two parabolas are defined by y=x^2-x-12 and y=-0.5x^2+5x+100. what is the distance between their vertices ?
Found 3 solutions by josgarithmetic, MathTherapy, Alan3354:
Answer by josgarithmetic(39617)

Their vertices are ( 1/2, -47/4)( 1/2, -49/4 ) and (5, 225/2 ).
Distance between them with Distance Formula...
..


124.33 124.846

Answer by MathTherapy(10551)


two parabolas are defined by y=x^2-x-12 and y=-0.5x^2+5x+100. what is the distance between their vertices ?
WRONG VERTEX!! From you know who! 


Answer by Alan3354(69443)

two parabolas are defined by y=x^2-x-12 and y=-0.5x^2+5x+100. what is the distance between their vertices ?
--------------
y=x^2-x-12
Vertex at (0.5,-12.25)
------------
y=-0.5x^2+5x+100
Vertex at (5,112.5)
============
distance = sqrt(4.5^2 + 124.75^2)
=~ 124.83 units